Areva Resources Namibia Plc. undertook a project involving the construction of a seawater desalination plant at Wlotzkasbaken near Swakopmund. The project was completed on 01 January 2017 at a cost of N$10 000 000. The directors of Areva Resources believe their multi-million dollar project shall be able to supply all the water to be consumed at Trikkopje mine, some 40km into the dessert. After 5 years, the company has an obligation to dismantle and restore the environment in compliance with both the Ministry of Marine Resources & Fisheries and the Ministry of Environment & Tourism regulations.
Management of Areva Resources has established that the cost of dismantling and restoration of the environment was N$4 000 000 on 1 January 2017. The year-onyear inflation rate was 6% on 1 January 2017 and is unlikely to change. The company specific pre-tax risk adjusted discount rate is 10%.
Required:
Prepare the relevant journal entries from 1 January 2017 to 31 December 2018 arising from the transactions above. Assume that on 31 December 2019 there was an advance in technology which would save the final cost of the dismantling and restoration cost by N$500 000.        
Answer #1

Dismantling and Restoration cost. = $4000000

Less: Cost reduced due to Tech advancement. = $(500000)

Total cost to be incurred. = $3500000R

RateAfter Inflation 6% for 5 years{3500000*(1.06)5= $4683790

Amortized in 5 years = 4683790/5 = $936758 every year

Journal entries

1.)

Reserves & Surplus A/C dr. $4683790
To Provision for Restoration $4683790

Being provision for dismantling and Restoration created to be amortized in 5 years.

2.)

Provision for Restoration A/c Dr. $936758
To. Profit and Loss A/c(Restoration Expenses) $936758

Being amortization of future Restoration expenses